OpenStudy (anonymous):

how would I solve 3^x^3 = 9^x

OpenStudy (anonymous):

\[3^{x ^{3}} = 9^{x}\]

OpenStudy (anonymous):

you mean this?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

well \( \Large 9^{x} = 3 ^{2x}\) so \( \Large 3^{x ^{3}} = 3 ^{2x}\) so you can say 2x = x^3

OpenStudy (anonymous):

I'm sorry but I don't understand

OpenStudy (anonymous):

as 9 = 3^2 so 9^x is (3^2)^x = 3 ^ 2x and it's equal to 3^x^3 so 3^x^3 = 3^2x

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

nice ;)

OpenStudy (anonymous):

Can you help me with another one?